OpenStudy (anonymous):

sqrt 6 (2sqrt 3 - 4 sqrt 2)

OpenStudy (anonymous):

\[\sqrt{6}(2\sqrt{3}-4\sqrt{2)} you need \to distribute the \sqrt{6}\]

OpenStudy (anonymous):

\[2\sqrt{18}-4\sqrt{12}\]

OpenStudy (anonymous):

take out 9 as a perfect square and 4 as a perfect square

OpenStudy (anonymous):

\[2\sqrt{9*2} - 4\sqrt{4*3}\]

OpenStudy (anonymous):

2*3\[2*3\sqrt{2} - 4*2\sqrt{3}\]

OpenStudy (anonymous):

\[6\sqrt{2}-8\sqrt{3}\]

OpenStudy (anonymous):

do you understand that

OpenStudy (anonymous):

yes thank you